OpenStudy (anonymous):

How does the volume of a sphere change if its radius doubles? (8th grade geometry)

OpenStudy (anonymous):

Thanks

Mehek (mehek14):

your welcome and close this question

OpenStudy (irishboy123):

V2 = (4/3) π (2R) ^3 V1 = (4/3) π (R) ^3